A diving bell is to be designed to withstand the pressure of seawater at a depth of

900 m.


1
Expert's answer
2021-09-02T10:41:25-0400

The pressure under water is given as follows:


"p = \rho g h"

where "h = 900m" is the depth, "g\approx 9.81m/s^2" is the gravitational acceleration, "\rho = 1025kg/m^3" is the density of seawater. Thus, obtain:


"p = 1025kg/m^3\cdot 9.81m/s^2\cdot 900m \approx 9.05\times 10^6Pa"

Answer. "9.05\times 10^6Pa".